how to work out and what is the LCM and HCF of 24 and 22 please answer ASAP
Answer by josgarithmetic(39800)

LCM is just like finding a common denominator. HCF, if Highest Common Factor, is like so:

24=2*12=2*2*2*3.
22=2*11.
Which factor occurs in BOTH numbers? Although "2" occurs in both numbers, it only occurs in 22, one time, and therefore 2 is the Highest Common Factor.

You can use prime factorization to find the LCM for 24 and 22. This is like raising fractions to higher terms...
The 2*2*2*3 is lacking 11, as found in 22.
The 2*11 is lacking in 2*2*3, as found in 24.
Can you figure the rest of this to obtain the Lowest Common Multiple of 22 and 24?